# Conformal Convolution and Monte Carlo Meta-learners for Predictive Inference of Individual Treatment Effects

### v4

@@ B.1 T-learner
1:  Input: data $X$, target $y$, treatment $W$, predictive distributions $Q_{Y^{0}}$ and $Q_{Y^{1}}$, MC samples $n$
2:  $ITE_{MC}=zeros(len(X)\cdot n)$
3:  for $i=1$ to $len(X)$ do
@@
6:           Sample $u_{1}\sim uniform(0,1,n$ times)
7:           $ITE_{MC,i\cdot n:(i+1)n}$ = $\textbf{Q}^{-1}_{Y^{1}}(X[i],u_{1})-y[i]$
@@
9:           Sample $u_{0}\sim uniform(0,1,n$ times)
10:           $ITE_{MC,i\cdot n:(i+1)n}$ = $y[i]$ $\textbf{Q}^{-1}_{Y^{0}}(X[i]$, $u_{0})$
@@
13:        Sample $u_{0}\sim uniform(0,1,n$ times)
14:        Sample $u_{1}\sim uniform(0,1,n$ times)
15:        $ITE_{MC,i\cdot n:(i+1)*n}$ = $\textbf{Q}^{-1}_{Y^{1}}(X[i],u_{1})-\textbf{Q}^{-1}_{Y^{0}}(X[i],u_{0})$
@@
B.2 S-Learner
The CMC S-learner closely resembles the T-learner algorithm with a notable distinction: it employs an S-learner as a CATE estimator and a single regression model for modeling the nuisance parameters. Consequently, we can only use split WCPS (SWCPS). This approach offers the advantage of leveraging a single model, thereby increasing the availability of samples to learn common components of the outcome. However, a drawback with this method arises from shared epistemic uncertainty, accounted for twice, potentially compromising the validity of the ITE predictive distributions. Nonetheless, the method remains capable of furnishing conservatively valid prediction intervals even in this scenario, as we observe in the simulation experiments.
B.3 X-Learner
@@
Appendix C Literature Overview
This section provides a more comprehensive overview of the literature on predictive inference ITE and quantifying related predictive uncertainty. Uncertainty quantification is a crucial component for treatment effect estimations for reliable decision-making in high-risk environments, such as healthcare [8]. In the past, most proposals for uncertainty quantification in ITE estimation used Bayesian approaches, such as Bayesian additive regression trees (BART) [10] and Gaussian processes [11]. The problem with these Bayesian approaches is that they are often uncalibrated without knowing the underlying data-generating process. These Bayesian approaches are also model-specific and are not easily transferred to modern ML approaches.
Recently, some works have tried to resolve this issue by using the CP framework  [12, 15], which can provide non-asymptotic distribution-free coverage guarantees and thus resolves the significant shortcomings of these Bayesian approaches. Lei and Candès [12] was the first work to propose a conformal inference approach to produce reliable prediction intervals for ITEs. More specifically, they proposed using WCP to account for covariate shifts as the distributions of covariates for treated and untreated subjects can differ from the target population. They proposed three variants of WCP ITE algorithms: (1) Naive WCP, which combines two potential outcome interval estimates, estimated by applying both WCP and a Bonferroni correction to get an ITE prediction interval; (2) Exact nested WCP, which first generates plug-in estimates of ITEs by using WCP and uses these estimates to apply a regular CP; and (3) Inexact nested WCP which replaces the second CP step in Exact WCP with quantile regression which does not provide any coverage guarantees. A significant drawback of the exact and inexact WCP approaches is that they require refitting the model, at least partially, for different confidence levels for the prediction intervals. Additionally, all the proposed approaches based on WCP only provide intervals as uncertainty quantification.
Alaa et al. [15] propose conformal meta (CM)-learners, which is a model-agnostic framework for issuing predictive intervals for ITEs that applies the inductive CP on top of CATE meta-learners. These CATE meta-learners are model-agnostic approaches that estimate the CATE by combining nuisance estimates and incorporating some inductive biases. These CM-learners are applied to a particular class of meta-learners, direct learners, which use a two-step procedure where nuisance estimators are fitted on the outcome in the first step. In the second step, these nuisance estimators are used to generate a pseudo-outcome in which the covariates are regressed on to get point estimates of the CATE; examples of such learners are X-, DR-, and IPW-learners. To construct prediction intervals of the ITE, Alaa et al. [15] propose to use these pseudo outcomes to generate conformity scores together with the estimates of the meta-learner, i.e., CATE estimator. However, their proposed framework requires more stringent assumptions to be valid than the classical ones for estimating CATE. They developed the so-called stochastic ordering framework and showed that CM-learners are marginally valid if their pseudo-outcome conformity scores are stochastically larger than the "oracle" conformity scores evaluated on the unobserved ITEs. However, it is difficult to translate this into practice, as the only way to verify this assumption is to observe the true ITEs, which is impossible. Additionally, their method is limited to only producing prediction intervals of ITE.
Appendix D Experimental setups
D.1 Synthetic Experiments
D.1.1 Alaa Synthetic Data Experiments

### v5

@@ Appendix C Conformal Monte Carlo meta-learners
1:  Input: observations $\{Z_{i}=(Y_{i},X_{i},W_{i})\}_{i=1}^{n}$, predictive distributions $Q_{Y(0)}$ and $Q_{Y(1)}$, MC samples $n_{MC}$
2:  $ITE_{MC}=zeros(n\cdot n_{MC})$
3:  for $i=1$ to $n$ do
@@
6:           Sample $u_{1}\sim Uniform(0,1,n_{MC}\text{ times})$
7:           $ITE_{MC,i\cdot n_{MC}:(i+1)n_{MC}}$ = $\textbf{Q}^{-1}_{Y(1)}(X_{i},u_{1},w_{1})-Y_{i}$where $Q^{-1}(x,u,w)=\inf\{y:Q(x,y,w)\geq u\}$ and uses weights specified in Table 3}
@@
9:           Sample $u_{0}\sim Uniform(0,1,n_{MC}\text{ times})$
10:           $ITE_{MC,i\cdot n_{MC}:(i+1)n_{MC}}=y_{i}-\textbf{Q}^{-1}_{Y(0)}(X_{i},u_{0},w_{0})$
@@
13:        Sample $u_{0}\sim Uniform(0,1,n_{MC}\text{ times})$
14:        Sample $u_{1}\sim Uniform(0,1,n_{MC}\text{ times})$
15:        $ITE_{MC,i\cdot n_{MC}:(i+1)n_{MC}}=\textbf{Q}^{-1}_{Y(1)}(X_{i},u_{1},w_{1})-\textbf{Q}^{-1}_{Y(0)}(X_{i},u_{0},w_{0})$
@@
C.1 S-learner
The CMC S-learner closely resembles the T-learner algorithm with a notable distinction: it employs an S-learner as a CATE estimator and a single regression model for modeling the nuisance parameters. This approach offers the advantage of leveraging a single model, thereby increasing the availability of samples to learn common components of the outcome. However, a drawback with this method arises from shared epistemic uncertainty, accounted for twice, potentially compromising the validity of the ITE predictive distributions. Nonetheless, the method remains capable of furnishing conservatively valid prediction intervals even in this scenario, as we observe in the simulation experiments.
C.2 X-learner
@@
Appendix D The Epsilon Problem: (un)identifiability of ITE predictive distribution
Beyond introducing the CCT and CMC meta-learners, we identify a fundamental challenge for predictive inference in ITE estimation: the unknown relationship between the noise distributions under treatment, $\varepsilon(1)$, and control, $\varepsilon(0)$. We propose a fourth assumption, in addition to the standard three outlined in Section 2, concerning the joint distribution of these noise terms. This assumption is critical for generating valid and calibrated predictive distributions within our framework.
Since $\varepsilon(0)$ and $\varepsilon(1)$ are never jointly observed, their joint distribution is not identifiable from data. Thus, any distributional regression approach requires an assumption about how these noise terms relate. In practice, most CATE literature assumes the noise terms are zero-mean, independent of covariates $(X,W)$, and either independent or identical across treatment arms [Künzel et al., 2019, Nie and Wager, 2021, Kennedy, 2022, Curth and Schaar, 2021]. Simulation studies often further assume either deterministic treatment effects (implying $\varepsilon(0)=\varepsilon(1)$) [Nie and Wager, 2021, Kennedy, 2022, Curth and Schaar, 2021] or full independence of the two noise terms [Künzel et al., 2019]. The limited literature concerning predictive inference of ITE  Hill [2011], Alaa and van der Schaar [2017], Lei and Candès [2021], Alaa et al. [2023], Vanderschueren et al. [2023] mostly assumes that the noise distributions are fully independent and sets up their experiments in the same way.
We formally state the identifiability constraint:
Proposition 1.
The ITE distribution $F_{Y(1)-Y(0)|X}(y)$ is only identifiable if
$F_{Y(0)|X}(y)\perp\!\!\!\perp F_{Y(1)|X}(y)$ (9)
or if a specific copula linking $Y(0)$ and $Y(1)$ is assumed.
Proof.
By definition,
[…]
Thus, the CDF of the difference depends on the full joint distribution $F_{Y(1),Y(0)}(y_{1},y_{0})$, not just on the two marginals.
Sklar’s theorem [Sklar, 1959] tells us that any joint CDF can be written in terms of its marginals and a copula $C$:
$F_{Y(1),Y(0)|X}(y_{1},y_{0})=C\left(F_{Y(1)|X}(y_{1}),F_{Y(0)|X}(y_{0})\right).$ (11)
Hence the difference CDF is
[…]
Thus, without specifying $C$, this integral, and hence $F_{Y(1)-Y(0)|X}(y)$, is not determined by the marginals alone.
Next, we present two cases where the ITE distribution is identifiable and one where it is not.
Case 1. Independence copula: If we assume $Y(0)\perp\!\!\!\perp Y(1)|X$, then $C(u_{1},u_{0})$, and thus,
[…]
which reduces to a convolution of two known marginals (where one is negated), and is therefore identifiable.
Case 2. Know copula: Generally, if one assumes a particulate copula $C$, then the joint law is specified by the two marginals and $C$, so the difference CDF becomes identifiable via the same integral above.
Case 3. No identified copula If neither the independence nor a known copula is assumed, then infinitely many different copulas are compatible with the same pair of marginals but induce different distributions of the difference $Y(1)-Y(0)$.
Therefore, no unique $F_{Y(1)-Y(0)|X}$ can be recovered from the marginals alone unless one imposes either the independence assumption or specifies the copula linking $Y(0)$ and $Y(1)$. ∎
Notably, the assumption of noise independence is already encoded in formal causal models like FFRCISTG [Robins, 1986] and Pearl’s nonparametric structural equation models.
Melnychuk et al. [2024] recently argue that ITE distributions are fundamentally unidentifiable in the potential outcomes framework, proposing partial identification via bounds on the CDF. While this is a promising direction, we take a more pragmatic stance: we assume a specific copula, namely the independent copula, for $\varepsilon(0)$ and $\varepsilon(1)$. Although this assumption is ultimately untestable and somewhat philosophical, we believe it provides a useful starting point for further research and discussion within the community.
D.1 Prediction intervals and violations of the noise assumption
In the following paragraphs, we will discuss three possible characteristics of the joint distribution of the noise terms and their impact on the validity of predictive intervals and ITE distributions when assuming independence. While the true nature of these joint noise distributions is, in part, a philosophical question, it requires further attention from the research community and domain experts.
$\mathbf{\boldsymbol{\varepsilon(0)=\varepsilon(1)}}$: Here, the ITE becomes deterministic, as the shared noise cancels out. This is often assumed in synthetic benchmarks. However, both CCT and CMC models will overestimate uncertainty, since they redundantly model shared aleatoric variability. A better approach would model epistemic uncertainty in both potential outcome models independently, allowing for more accurate uncertainty propagation via convolution or Monte Carlo sampling.
[…]
Simulation results (Figure 11) empirically validate these scenarios.
Appendix E Proofs
E.1 Probabilistic calibration of weighted conformal predictive system
Definition 1 (Tibshirani et al. [2019], Definition 1).
Random variables $S_{1:n}$ are said to be weighted exchangeable, with functions $w_{1:n}$, if the density $f$ of their joint distribution can be factorized as
$f(s_{1},\ldots,s_{n})=\prod_{i=1}^{n}w_{i}(s_{i})g(s_{1},\ldots,s_{n}),$ (14)
where $g$ is any function that does not depend on the ordering of its inputs, i.e., $g(s_{\sigma(1)},\ldots,s_{\sigma(n)})$ for any permutation $\sigma$ of $1,\ldots,n$.
See 1
Proof.
The proof strategy is inspired by the proof of smoothed conformal prediction (Angelopoulos et al. [2024], Lemma 9.3).
Let $E_{z}$ denote the event that $\{Z_{1},\ldots,Z_{n+1}\}=\{z_{1},\ldots,z_{n+1}\}$, and let the $s_{i}=s(z_{i};z_{1:n+1\setminus i})$ for $i=1,\ldots,n+1$. Let $f$ be the density function of the joint sample $Z_{1:n+1}$. For each $i$,
$\mathbb{P}\left\{S_{n+1}=s_{i}|E_{z}\right\}=\mathbb{P}\left\{Z_{n+1}=z_{i}|E_{z}\right\}=\frac{\sum_{\sigma|\sigma(n+1)=i}f(z_{\sigma(1):\sigma(n+1)})}{\sum_{\sigma}f(z_{\sigma(1):\sigma(n+1)})}.$
Since $Z_{1:n+1}$ are weighted exchangeable (Definition 1),
$\begin{split}\frac{\sum_{\sigma|\sigma(n+1)=i}f(z_{\sigma(1):\sigma(n+1)})}{\sum_{\sigma}f(z_{\sigma(1):\sigma(n+1)})}&=\frac{\sum_{\sigma|\sigma(n+1)=i}\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})g(z_{\sigma(1):\sigma(n+1)})}{\sum_{\sigma}\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})g(z_{\sigma(1):\sigma(n+1)})}\\ &=\frac{\sum_{\sigma|\sigma(n+1)=i}\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})g(z_{1:n+1})}{\sum_{\sigma}\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})g(z_{1:n+1})}\\ &=\frac{\sum_{\sigma|\sigma(n+1)=i}\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})}{\sum_{\sigma}\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})}\\ &\stackrel{{\scriptstyle(*)}}{{=}}\frac{w(x_{i})}{\sum_{j=1}^{n+1}w(x_{j})}=p_{i}^{w}\end{split}$ (15)
The simplification at ()(*)( ∗ ) can be made because we consider only a special case of weighted exchangeability, the covariate shift, where $w_{1:n}\equiv 1$ and $w_{n+1}((x,y))=w(x)=\frac{d\tilde{P}_{X}(x)}{dP(x)}$. Nevertheless, our findings can be readily applied to the general concept of weighted exchangeability.
The result in 15 shows that,
$S_{n+1}|E_{z}\sim\sum_{i=1}^{n+1}p_{i}^{w}\delta_{s_{i}}.$ (16)
For each $j\in[1:n+1]$ define,
$p_{j}(\phi)=\sum_{i=1}^{n+1}[S_{i}<S_{j}^{y}]p_{i}^{w}(X_{j})+\sum_{i=1}^{n+1}[S_{i}^{y}=S_{j}^{y}]p_{i}^{w}(X_{j})\phi.$ (17)
where $S_{i}=S_{i}^{Y_{n+1}}$ for any $i\in[1:n+1]$.
By construction, the result in 16, and the tower law,
$\mathbb{P}(p_{n+1}(\phi)\leq\alpha)=\sum_{j=1}^{n+1}p_{j}^{w}\mathbb{P}(p_{j}(\phi)\leq\alpha)=\mathbb{E}\left[\sum_{j=1}^{n+1}p_{j}^{w}\mathbb{P}(p_{j}(\phi)\leq\alpha|S_{1:n+1})\right].$ (18)
Define the weighted quantile,
$q=\text{Quantile}\left(\sum_{i=1}^{n+1}p_{i}^{w}(X_{i})\delta_{S_{i}};1-\alpha\right),$
and let
$N^{-}=\sum_{j=1}^{n+1}[S_{j}<q]p_{j}^{w},\qquad N^{=}=\sum_{j=1}^{n+1}[S_{j}=q]p_{j}^{w}.$
By definition of the weighted quantile, for any $j$ if $S_{j}=q$,
$p_{j}({\phi})=N^{-}+\phi N^{=},$
and thus,
$p_{j}({\phi})\leq\alpha\iff\phi\leq\frac{\alpha-N^{-}}{N^{=}}$
Therefore,
$\mathbb{P}(p_{j}(\phi)\leq\alpha|S_{1:n+1})=[S_{j}<q]+[S_{j}=q]\frac{\alpha-N^{-}}{N^{=}}.$ (19)
Hence, combined with the results of 18,
$\begin{split}\mathbb{P}(p_{j}(\phi)\leq\alpha)&=\mathbb{E}\left[\sum_{j=1}^{n+1}p_{j}^{w}\mathbb{P}(p_{j}(\phi)\leq\alpha|S_{1:n+1})\right]\\ &=\mathbb{E}\left[\sum_{j=1}^{n+1}p_{j}^{w}\left\{[S_{j}<q]+[S_{j}=q]\frac{\alpha-N^{-}}{N^{=}}\right\}\right]\\ &=\mathbb{E}\left[\sum_{j=1}^{n+1}[S_{j}<q]p_{j}^{w}+\sum_{j=1}^{n+1}[S_{j}=q]p_{j}^{w}\frac{\alpha-N^{-}}{N^{=}}\right]\\ &=\mathbb{E}\left[N^{-}+N^{=}\frac{\alpha-N^{-}}{N^{=}}\right]=\mathbb{E}\left[\alpha\right]=\alpha.\end{split}$ (20)
Thus, the distribution of $Q\left(X_{n+1},Y_{n+1},\phi,\frac{d\tilde{P}}{dP};Z_{1:n}\right)$ is uniform, completing the proof. ∎
E.2 Probabilistic calibration of CCT-learner
See 2
Proof.
The proof follows directly from Theorem 1, where we assume weighted exchangeability according to,
$Y_{n+1}(1)-Y_{n+1}(0)|E_{q_{Y(1)}-q_{Y(0)}}\sim\sum_{i=1}^{n_{0}}\sum_{j=1}^{n_{1}}p^{w_{0}}_{i}p^{w_{j}}_{i}\delta_{q_{Y(1),j}-q_{Y(0),i}}+[1-\{1-p_{n+1}^{w_{1}}\}\{1-p_{n+1}^{w_{0}}\}]\delta_{\infty}.$ (21)
E.3 Existence universal consistent CCT- and CMC T-learner
The following section presents and proves Theorem 3, which states the existence of a universal consistent CCT- and CMC T-learner, i.e., the conformal predictive distribution approaches the true data generating distribution as $n\rightarrow\infty$ under any data-generating distribution. We introduce here Definition 2, taken in verbatim from Vovk [2022], which lays out our notion of consistency for CPS.
Definition 2 (in verbatim Vovk [2022]).
A conformal predictive system $Q$ is consistent for a probability measure $\mathcal{P}$ on $Z:=X\times\mathbb{R}$ if, for any bounded continuous function $f:\mathbb{R}\rightarrow\mathbb{R}$,
[…]
in probability, where:
  • $Q_{n}$ is the predictive distribution $Q_{n}:y\mapsto Q(X_{n+1},\phi;Z1:n)$ output by Q as its forecast for the label $Y_{n+1}$ corresponding to the test object $X_{n+1}$ based on the training data $Z_{1:n}$, where $Z_{i}=(X_{i},Y_{i})$ for all $i$.
  • $\mathbb{E}_{\mathcal{P}}(f|X_{n+1})$ is the conditional expectation of $f(y)$ given $x=X_{n+1}$ under $(x,y)\sim\mathcal{P}$.
  • $Z_{i}=(X_{i},Y_{i})\sim\mathcal{P},i=1,...,n$, $Z_{n+1}\sim\mathcal{P}$, and $\phi\sim Uniform(0,1)$, are assumed all independent.
See 3
Proof.
The philosophy behind the MC simulations in the CMC T-learner is to not perform a convolution of both outcome predictive distributions, as this would quickly become an expensive operation that would need to be performed for every example during inference. However, to prove that there exists a universal consistent learner, we replace this MC sampling in Algorithm 2 with the convolution operation,
$\begin{split}Q_{ITE}&(X_{n+1},y,\phi,\pi;Z_{1:n})=\\ &\int Q_{Y(1)}(X_{n+1},y+t,\phi,w_{1}=\frac{1}{\pi};Z_{1:n1}^{W=1})dQ_{Y(0)}(X_{n+1},t,\phi,\frac{1}{1-\pi};Z_{1:n0}^{W=0}).\end{split}$ (23)
Thus, becoming a CCT-learner is equivalent to the CMC T-learner approach asymptotically.
Considering the existence of a universal consistent learner, we no longer need to worry about the covariate shift. We can replace the WCPS with CPS in the CCT- and CMC T-learner. This is equivalent to setting the propensity score to $0.5$ for every covariate.
By Theorem 3 in Vovk [2022], we know that there exists universal CPS if the object space $\mathcal{X}$ is standard Borel, meaning that for any probability measure $\mathcal{P}$ on $Z$, the CPS is consistent. Consequently, a universal CPS exists that can model the true data-generating distribution outcome under treatment $Y(1)$ and control $Y(0)$. These distributions are weakly converging distributions. Since we assume that the predictive distributions of the outcome are independent,
$Y(0)\perp\!\!\!\perp Y(1)|X,$ (24)
the outcome variables are independent conditionally on the covariate set $X$. Then, by the definition of convolution and the fact that $ITE=Y(1)-Y(0)$, the convolution of the true conditional cumulative distribution of outcome under treatment $F_{Y(1)|X}$ and outcome under negative control $F_{-Y(0)|X}$ equals the true conditional cumulative ITE distribution,
[…]
Appendix F Experiments results
Figure 3: Simulation results for synthetic experiment setups A and B from Alaa et al. [2023].
Figure 4: Simulation results for synthetic experiment setup A Nie and Wager [2021].
Figure 5: Simulation results for synthetic experiment setup B Nie and Wager [2021].
Figure 6: Simulation results for synthetic experiment setup C Nie and Wager [2021].
Figure 7: Simulation results for synthetic experiment setup D Nie and Wager [2021].
Figure 8: Simulation results for semi-synthetic experiments EDU and IHDP.
Figure 9: Observed frequency of PIT-values on the synthetic experiments of Alaa et al. [2023] and the semi-synthetic experiment EDU.
F.1 Conformal prediction intervals
Both CCT- and CMC meta-learners return conditional predictive distributions of the ITE. Prediction intervals of confidence level $1-\alpha$ are formed by taking the $\frac{\alpha}{2}$ and $1-\frac{\alpha}{2}$ quantiles from the predictive distribution as the lower and upper bounds, respectively. Our method produces symmetric, two-sided intervals with equally distributed probability density. In contrast, the approaches by Lei and Candès [2021] and Alaa et al. [2023] generate asymmetric intervals, offering an efficiency advantage due to smaller interval lengths [Romano et al., 2019]. However, these asymmetric intervals can be misleading in practical applications as they might intuitively imply equal probability density on both sides of the interval.
The experiments assessed coverage and efficiency of prediction intervals of the CCT-learner and (pseudo) CMC T-learner against WCP (exact, inexact, and naive) [Lei and Candès, 2021], CM-Learner (DR) [Alaa et al., 2023], and a CPS Oracle fitted as a standard regression problem on the ITE.
Figure 10: Simulation results for setups A and B from Alaa et al. [2023] for 90% ITE prediction interval.
Coverage and efficiency of prediction intervals
On synthetic datasets, the CCT- and CMC meta-learners consistently demonstrated slightly conservative coverage across various simulations while maintaining high efficiency, indicated by small average interval widths. In contrast, the inexact WCP, while more efficient in some cases, often did so at the cost of significant under coverage. The CM DR learner was consistently more conservative than our proposed methods. Other WCP approaches, with proven conservative validity, were extremely conservative across all confidence levels and experiments, likely due to their over-conservative design, employing either Bonferroni correction or double conformal inference procedure [Lei and Candès, 2021].
For the three semi-synthetic datasets, we performed experiments with covariates from real-world data. As here no propensity scores are provided, these resemble a realistic scenario. We estimate the propensity scores using logistic regression. The most extensive experiment was conducted on the 77 settings of the ACIC2016 dataset Dorie et al. [2019], which contain 58 covariates (3 categorical, 5 binary, 27 counts, 23 continuous) with varying complexity of outcome surfaces, degree of confounding, overlap, and treatment effect heterogeneity. Results, depicted in Figure 12, show that our proposed approaches provide conservatively valid prediction intervals across all settings, with significant conservative coverage in most settings, though less than the CM DR-learner and valid WCP approaches, likely due to violations in the assumption of conditional independence of the noise distributions. Regarding efficiency, the CCT-learner performed best for the reasons previously mentioned. As can be seen in Figure 17, a similar conclusion can be drawn for the other semi-synthetic experiments.
Performance under different relations between noise distributions
Appendix G Experimental setups
G.1 Experiments compute resources
All experiments were conducted on a private compute cluster. Each experimental run utilized 4 CPU cores and 30 GB of memory, with no GPU acceleration required. For experiments involving random forest-based methods, individual runs completed in under 30 seconds, while other configurations (e.g., preprocessing, hyperparameter tuning) exhibited longer execution times proportional to dataset size and algorithmic complexity.
G.2 Synthetic experiments
G.2.1 Alaa et al. synthetic data experiments